# ElastoGen: 4D Generative Elastodynamics

### v2

@@ A.1 Supplemental video
A.2 Diffusion network $\mathcal{D}$
The goal is to train a diffusion network $\mathcal{D}$ to generate the weights $\mathbf{W}$ of a corresponding NeuralMTL model $\mathcal{N}$, given the material parameters $\{e,\nu\}$. Here, $\mathbf{W}$ denotes the weights of $\mathcal{N}$, and the process is formulated as a conditional diffusion problem guided by $\{e,\nu\}$, such that $\mathbf{W}=\mathcal{D}(e,\nu)$.
To this end, we first construct a dataset consisting of 1000 paired samples of $\{e,\nu\}$ and $\mathbf{W}$, as described in § 4.3. Following the approach of Wang et al. (2024a), we utilize Latent Diffusion Models (LDM,  Rombach et al. (2022)) to generate $\mathbf{W}$, as our preliminary experiments showed that directly learning $\mathbf{W}$ led to suboptimal performance. To address this, we train an autoencoder to map the network weights $\mathbf{W}$ to a 256-dimensional latent vector, in which the diffusion process is performed.
When training the diffusion model, the autoencoder remains fixed, serving solely to encode $\mathbf{W}$ into its latent representation $l$. At each diffusion timestep $t$, we introduce noise $\epsilon_{t}$ to $l$, resulting in $l_{t}=l+\epsilon_{t}$. The objective is to train a noise prediction model, $\epsilon_{\theta}(l_{t},t;e,\nu)$, to estimate the noise $\epsilon_{t}$ at each timestep $t$, as described in § 3.2. During inference, we begin with random noise and progressively remove noise from it using the noise prediction model $\epsilon_{\theta}$, guided by the material parameters ${e,\nu}$. This iterative denoising process produces a 256-dimensional latent vector, which is subsequently passed through the decoder to generate the corresponding network weights $\mathbf{W}$.
We train the autoencoder using a learning rate of $1\times 10^{-3}$ and the diffusion model with a learning rate of $1\times 10^{-4}$. Both models are trained for 1000 epochs with a batch size of 64. The architecture of the autoencoder and diffusion model is detailed in table 2. Note that in diffusion process the 256-dimensional latent vector is viewed as a 1-channel $16\times 16$ image.
Network Layers ##\##Output features Description
@@ A.2 Diffusion network $\mathcal{D}$
A.3 Convolutional deformation gradient
Given an input 3D object, ElastoGen rasterizes it into a set of 3D cubes or voxels. For $i$-th sub-volume inside the 3D cubes, ElastoGen uses a 3D CNN to calculate $\mathcal{G}_{i}$. As $\mathcal{G}_{i}$ has an analytic format as described in equation 8, the kernel’s weights of 3D CNN can be directly computed. To be more clear, for $i$-th sub-volume containing 8 vertices, let $\mathbf{A}_{i}=[\mathbf{q}_{1},\mathbf{q}_{2},...\mathbf{q}_{8}]\in\mathbb{R}^{3\times 8}$ and $\bar{\mathbf{A}}_{i}=[\bar{\mathbf{q}}_{1},\bar{\mathbf{q}}_{2},...\bar{\mathbf{q}}_{8}]\in\mathbb{R}^{3\times 8}$ be deformed and rest-shape position of the vertices, the weights of 3D CNN can be filled with $\left[\big{(}\bar{\mathbf{A}}\bar{\mathbf{A}}^{\top}\big{)}^{-1}\bar{\mathbf{A}}\otimes\mathbf{I}\right]\in\mathbb{R}^{9\times 24}$. Here, the 3D CNN has an input channel of $3$, an output channel of $9$ and a kernel size of $2\times 2\times 2$.
A.4 Global phase
As stated in the main text, we need to solve the global linear system (equation 10), which requires determining $\mathbf{L}_{i}$ and $\mathbf{b}_{i}$. We abbreviate the neural strain $\mathcal{N}(\mathcal{G}_{i}[\mathbf{q}_{i}])$ as $\mathcal{N}$, rewriting equation 5, the energy $E_{i}$ for element $i$ is
@@ A.4 Global phase
Based on $\mathbf{L}_{i}\mathbf{q}-\mathbf{b}_{i}:=\frac{\partial E_{i}}{\partial\mathbf{q}}$, we can obtain the expression for $\mathbf{b}_{i}$ and $\mathbf{L}_{i}$. Taking the derivative of equation 11 with respective to position $\mathbf{q}$ we obtain
$\frac{\partial E_{i}}{\partial\mathbf{q}}=\omega_{i}\left(\mathbf{G}_{i}\mathcal{N}\mathcal{N}^{\top}\mathbf{G}_{i}^{\top}\mathbf{q}-\mathbf{U}_{i}\mathbf{V}_{i}^{\top}\mathcal{N}\mathbf{G}_{i}^{\top}\right),$ (12)
where $\mathbf{G}_{i}$ is $i$-th component of $\mathbf{G}$ corresponding to element $i$. Therefore, we derive $\mathbf{L}_{i}$ and $\mathbf{b}_{i}$ as
$\mathbf{L}_{i}=\omega_{i}\mathbf{G}_{i}\mathcal{N}\mathcal{N}^{\top}\mathbf{G}_{i}^{\top},\quad\mathbf{b}_{i}=\omega_{i}\mathbf{U}_{i}\mathbf{V}_{i}^{\top}\mathcal{N}\mathbf{G}_{i}^{\top}.$ (13)
As it indicates, for each voxel, we can obtain $\mathbf{b}_{i}$ by applying the transformation $\mathbf{G}_{i}^{\top}$ to $\mathbf{U}_{i}\mathbf{V}_{i}^{\top}\mathcal{N}$. For $\mathbf{G}_{i}$ has been trained as a convolutional kernel as described in § A.3, we can directly fetch the previously trained kernel and perform this operation.
For the linear system in equation 10, we further write it as $\mathbf{A}\mathbf{q}=\mathbf{b}$. For any diagonally dominant matrix $\mathbf{A}$, the linear system $\mathbf{A}\mathbf{q}=\mathbf{b}$ can be solved using iterative method as:
@@
where $\mathbf{D}$ is the diagonal part of $\mathbf{A}$ and the off-diagonal part $\mathbf{B}=\mathbf{A}-\mathbf{D}$, and $\mathbf{q}^{k}$ is the result after $k$ loops of RNN-2. In our case, $\mathbf{A}=\frac{\mathbf{M}}{h^{2}}+\sum_{i}\mathbf{L}_{i}$ and $\mathbf{b}=\mathbf{f}_{q}+\frac{\mathbf{M}}{h^{2}}\left(\mathbf{q}_{n}+h\dot{\mathbf{q}}_{n}\right)+\sum_{i}\mathbf{b}_{i}$ according to equation 10. Note that we use subscript $n$ to indicate timestep and superscript $k$ as index for RNN-2 loops.
Similar to § A.3, we use a 3D CNN to implement each iteration in RNN-2. The weights of 3D CNN are filled with $-\mathbf{D}^{-1}\mathbf{B}$ and the bias is filled with $\mathbf{D}^{-1}\mathbf{b}$. The number of input channels is $78$, and the number of output channels is $24$, with a kernel size of $1$, representing the contribution of each voxel to its $8$ vertices. The iterative process is formulated as a recurrent network, i.e. RNN-2 in our paper, to solve the global system.
A.5 Broader impact
Our model integrates computational physics knowledge into the network structure design, significantly reducing the data requirements and making both the training and network structure more lightweight. It blends the boundaries among machine learning, graphics, and computational physics, providing new perspectives for network design. Our model does not necessarily bring about any significant ethical considerations.
A.6 More quantitative validations
We compare the NeuralMTL strain with the ground truth under various deformed configurations. In each case, the neural energy models closely match the ground truth, demonstrating the effectiveness and expressiveness of our neural approximations for these nonlinear energy functions.
@@ A.6 More quantitative validations
A.7 More experiments
@@
We provide additional results in Fig. 10 and Fig. 11 to demonstrate the robustness of ElastoGen. For more animated results, we refer the readers to supplemental video.

### v3

@@ Appendix A Supplemental video
Appendix B Diffusion network
The goal is to train a diffusion network $\mathcal{D}$ to generate the weights $\mathbf{W}$ of a corresponding NeuralMTL $\mathcal{N}$, given the material parameters $\{e,\nu\}$. Here, $\mathbf{W}$ denotes the weights of $\mathcal{N}$, and the process is formulated as a conditional diffusion problem guided by $\{e,\nu\}$, such that $\mathbf{W}=\mathcal{D}(e,\nu)$.
To this end, we first construct a dataset consisting of 1000 paired samples of $\{e,\nu\}$ and $\mathbf{W}$, as described in § Decoupling NeuralMTL from deformation and material parameters. Following the approach of Wang et al(wang2024neural), we utilize latent diffusion models (LDM) (rombach2022high) to generate $\mathbf{W}$, as our preliminary experiments show that directly learning $\mathbf{W}$ leads to suboptimal performance. To address this, we train an autoencoder to map the network weights $\mathbf{W}$ to a 256-dimensional latent vector, in which the diffusion process is performed.
When training the diffusion model, the autoencoder remains fixed, serving solely to encode $\mathbf{W}$ into its latent representation $l$. At each diffusion timestep $t$, we introduce noise $\epsilon_{t}$ to $l$, resulting in $l_{t}=l+\epsilon_{t}$. The objective is to train a noise prediction model, $\epsilon_{\theta}(l_{t},t;e,\nu)$, to estimate the noise $\epsilon_{t}$ at each timestep $t$, as described in § Diffusion model. During inference, we begin with random noise and progressively remove noise from it using the noise prediction model $\epsilon_{\theta}$, guided by the material parameters $\{e,\nu\}$. This iterative denoising process produces a 256-dimensional latent vector, which is subsequently passed through the decoder to generate the corresponding network weights $\mathbf{W}$.
We train the autoencoder using a learning rate of $1\times 10^{-3}$ and the diffusion model with a learning rate of $1\times 10^{-4}$. Both models are trained for 1000 epochs with a batch size of 64. The architecture of the autoencoder and diffusion model is detailed in Tab. 3. Note that in diffusion process the 256-dimensional latent vector is viewed as a 1-channel $16\times 16$ image.
Network Layers #\#Output features Description
@@ Appendix B Diffusion network
Appendix C Convolutional deformation gradient
Given an input 3D object, ElastoGen rasterizes it into a set of 3D voxels. For $i$-th voxel, ElastoGen uses a 3D CNN to calculate $\mathcal{G}_{i}$. As $\mathcal{G}_{i}$ has an analytic format as described in Eq. (8), the kernel’s weights of 3D CNN can be directly computed. In details, for $i$-th voxel containing 8 vertices, let $\mathbf{A}_{i}=[\mathbf{q}_{1},\mathbf{q}_{2},...\mathbf{q}_{8}]\in\mathbb{R}^{3\times 8}$ and $\bar{\mathbf{A}}_{i}=[\bar{\mathbf{q}}_{1},\bar{\mathbf{q}}_{2},...\bar{\mathbf{q}}_{8}]\in\mathbb{R}^{3\times 8}$ be deformed and rest-shape position of the vertices respectively, the weights of 3D CNN can be filled with $\left[\big(\bar{\mathbf{A}}\bar{\mathbf{A}}^{\top}\big)^{-1}\bar{\mathbf{A}}\otimes\mathbf{I}\right]\in\mathbb{R}^{9\times 24}$. Here, the 3D CNN has an input channel of $3$, an output channel of $9$ and a kernel size of $2\times 2\times 2$.
Appendix D Global phase
As stated in the main text, we need to solve the global linear system as in Eq. (10), which requires determining $\mathbf{L}_{i}$ and $\mathbf{b}_{i}$. We abbreviate the neural strain $\mathcal{N}(\mathcal{G}_{i}[\mathbf{q}_{i}])$ as $\mathcal{N}$, rewriting Eq. (5), the energy $E_{i}$ for voxel $i$ is
@@ Appendix D Global phase
For the convenience of subsequent derivations, we rewrite Eq. (11) as:
[…]
where $\mathbf{N}=\mathcal{N}^{\top}\otimes\mathbf{I}$, $\mathrm{vec}(\cdot)$ flattens a matrix into a vector, $\mathbf{G}_{i}$ is a linear operator projects DOFs $\mathbf{q}$ to the i-th element’s deformation gradient $\mathbf{F}_{i}$, and $\mathbf{R}_{i}=\mathbf{U}_{i}\mathbf{V}_{i}^{\top}$. Taking the derivative of Eq. (12) with respect to position $\mathbf{q}$ we obtain
$\frac{\partial E_{i}}{\partial\mathbf{q}}=\omega_{i}\left(\mathbf{G}_{i}^{\top}\mathbf{N}^{\top}\mathbf{N}\mathbf{G}_{i}\mathbf{q}-\mathbf{G}_{i}^{\top}\mathbf{N}^{\top}\mathrm{vec}(\mathbf{R}_{i})\right).$ (13)
Comparing it to the definition, $\mathbf{L}_{i}\mathbf{q}-\mathbf{b}_{i}:=\frac{\partial E_{i}}{\partial\mathbf{q}}$, we obtain the expression for $\mathbf{b}_{i}$ and $\mathbf{L}_{i}$ as
$\mathbf{L}_{i}=\omega_{i}\mathbf{G}_{i}^{\top}\mathbf{N}^{\top}\mathbf{N}\mathbf{G}_{i},\quad\mathbf{b}_{i}=\omega_{i}\mathbf{G}_{i}^{\top}\mathbf{N}^{\top}\mathrm{vec}(\mathbf{R}_{i}).$ (14)
As it indicates, for each voxel, we can obtain $\mathbf{b}_{i}$ by applying the transformation $\mathbf{G}_{i}^{\top}$ to $\mathbf{N}^{\top}\mathrm{vec}(\mathbf{R}_{i})$. For $\mathbf{G}_{i}$ has been trained as a convolutional kernel as described in § C, we can directly fetch the previously trained kernel and perform transposed 3D convolution.
For the linear system in Eq. (10), we rewrite it as $\mathbf{A}\mathbf{q}=\mathbf{b}$ for brevity. For any diagonally dominant matrix $\mathbf{A}$, the linear system $\mathbf{A}\mathbf{q}=\mathbf{b}$ can be solved using iterative method as:
@@
where $\mathbf{D}$ is the diagonal part of $\mathbf{A}$ and the off-diagonal part $\mathbf{B}=\mathbf{A}-\mathbf{D}$, and $\mathbf{q}^{k}$ is the result after $k$ loops of RNN-2. In our case, $\mathbf{A}=\frac{\mathbf{M}}{h^{2}}+\sum_{i}\mathbf{L}_{i}$ and $\mathbf{b}=\mathbf{f}_{q}+\frac{\mathbf{M}}{h^{2}}\left(\mathbf{q}^{n}+h\dot{\mathbf{q}}^{n}\right)+\sum_{i}\mathbf{b}_{i}$ according to Eq. (10). Note that we use superscript $n$ to indicate timestep and superscript $k$ as the index for RNN-2 loops.
@@
Similar to § C, to compute $\sum_{i}\mathbf{L}_{i}\mathbf{q}^{k}$, we first apply $\mathbf{G}_{i}$ to $\mathbf{q}^{k}$, then right-multiply the result with $\omega_{i}\mathbf{N}^{\top}\mathbf{N}$, and finally apply $\mathbf{G}_{i}^{\top}$ for each voxel. This process can be implemented using a 3D contolution, a matrix multiplication, and a transposed 3D convolution. The iterative process is formulated as a recurrent network (i.e., RNN-2) to solve the global system.
Appendix E Broader impact
Our model integrates computational physics knowledge into the network structure design, significantly reducing the data requirements and making both the training and network structure more lightweight. It blends the boundaries of machine learning, graphics, and computational physics, providing new perspectives for network design. Our model does not necessarily bring about any significant ethical considerations.
Appendix F More quantitative validations
We compare the NeuralMTL strain with the ground truth under various deformed configurations as in Fig. 8. In each case, the neural energy models closely match the ground truth, demonstrating the effectiveness and expressiveness of our neural approximations for these nonlinear energy functions.
Appendix G Convergence study
To quantify the impact of RNN loops and the subspace encoding, we compare ElastoGen predictions using different RNN loops with the ground truth, computed via solving the global matrix with a direct solver, in terms of relative error. Specifically, the variational formulation requires that the derivative of the total energy (elastic + kinetic – external work) vanishes at equilibrium. Since the total derivative equals zero at convergence, we define the ground-truth residual as the negative of known external forces (e.g., gravity, boundary forces), and compare it to the predicted residual composed of the derivatives of elastic and kinetic energies.
The results and convergence plots are shown in Fig. 9. In this test, one end of the beam is fixed, and ElastoGen predicts its twisting trajectory under forces. We observe that 50 RNN iterations converge the ElastoGen prediction to GT. Reducing the loop count to 20 still yields satisfactory results, while using only 1, 3, or 5 iterations leads to noticeably stiffer dynamics. In this experiment, RNN-2 employs an 18-dimensional subspace encoder to extract low-frequency residuals. Without this encoding, local relaxation fails to converge.
Appendix H More experiments
@@
We provide additional results in Fig. 10 and Fig. 11 to demonstrate the robustness of ElastoGen. For animated results, we refer the readers to supplemental video.